OpenStudy (fluttershyk):

8 to the 2 power +9(12/ 3*2)-7

OpenStudy (fluttershyk):

@rebeccaxhawaii @KamiBug

OpenStudy (fluttershyk):

I know the answer is 129 I just need to know the steps

OpenStudy (fluttershyk):

@pooja195

jhonyy9 (jhonyy9):

what are there please exactly ? 8^2 + 9(12/(3*2)) -7 = ? this is right wrote ?

jhonyy9 (jhonyy9):

inside parentheses are 12 divide 3 time 2 so 12 divide 6 what is equal 2 ?

jhonyy9 (jhonyy9):

please collaborate

jhonyy9 (jhonyy9):

i can help you sure

OpenStudy (fluttershyk):

8^2 + 9(12/(3*2)) -7 = ? this is written write I want to know the steps to solve it I know the answer is 129 I just need to know the steps to get 129

jhonyy9 (jhonyy9):

ok. the first step will be 3*2 = 6 bc. these are inside parentheses inside parentheses so the second step 12/6 3rd this result multiply by 9 and make the second power the 8 assume to this result from parentheses multiplied by 9 and subtract from this the 7

jhonyy9 (jhonyy9):

hope helped

jhonyy9 (jhonyy9):

but making these what i ve wrote above dont will result to you 129 never i think that inside parentheses need being (12/3)*2 = 8 so 9*8=72-7=65 and 64 +65 = 129 this is the way to get 129 hope this is understandably sure right easy now

jhonyy9 (jhonyy9):

ok. ?

OpenStudy (fluttershyk):

I got it

jhonyy9 (jhonyy9):

was my pleasure good luck and anytime bye bye